sir there r three ques that i m not getting jst a simple concept in them plz help with complete solution
1)find the point on y axis which is equidistance from points(5,5,2) & (3,1,2)

2)find tha points on tha x-y plane which is equidistance from points (0,3,2)
(2,0,3) and (0,0,1)

3)a point R with z- coordinate 8 lies on the line segment joining the points
p(2,-3,4) &q(8,0,10)
find the coordinates of R.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear cool gal

1)   let a point on y axix (0,y,0)

now use distance formula

[(5-0)2 +(5-y)2 +(2-0)2]1/2  = [(3-0)2 +(1-y)2 +(2-0)2]1/2

or

[(5-0)2 +(5-y)2 +(2-0)2]  = [(3-0)2 +(1-y)2 +(2-0)2]

or

25 + (5-y)=9+ (1-y)2

16 =(1-y)2 - (5-y)

16 =(1-y-5+y)(1-y+5-y)

16 =-4(6-2y)

40 = 8y

y=5


so point is  (0,5,0)


plese post different question in different post.

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin